OpenStudy (anonymous):

Simplify. Express the product as a radical expression. \[\sqrt[11]{x^5} \times \sqrt{x}\]

OpenStudy (anonymous):

i think its \[\sqrt[22]{x^5}\]

OpenStudy (anonymous):

\[\sqrt[11]{x^5}= x^{5/11}\] \[\sqrt{x}=x^{1/2}\] Now, \[\sqrt[11]{x^{5}}*\sqrt{x} = x^{5/11}*x^{1/2}\] \[= x^{5/11 + 1/2}\] \[=x^{21/22}\] or \[=\sqrt[22]{x^{21}}\]